The equation for the deflection along a particular uniform beam under a given load is given by da y cos(x) H(x – 2n7) = withn=9
Answer #1
  • EH1n-2nn) (-uosn) + Hsinn - this - antiH dn2 Integration again white +fnisine-fonde - dy - (H/n-an) (-won) dut dn fannede dy4 dy = Hn-2nn) (-sinn) -24cosh - H1² - 200 H x + (2+ 6+81) dn ď Integration y (41o =åt») ksion)-fantes da -Impedan - Jannten
  • final y = H (u-2017) issn - 34 sinn - Hus - noun ² + (2+ 6487²) Hn 4 120835,147H​​​​​​​